Études Anglaises, vol. 52, no. 3, 1999, p. 275–285

ARTISTIC REBIRTH AND THE IMAGES OF ANDROGYNY IN WILLA CATHER’S THE SONG OF THE LARK

The Song of the Lark, Willa Cather’s 1915 Künstlerroman set in the two decades that span the nineteenth and twentieth centuries, takes a second-generation Scandinavian immigrant girl from her native desert town of Moonstone, Colorado, to the great operas of the Western World. In the tradition of the Bildungsromane and of their male heroes, Cather has her heroine, Thea Kronborg, go the hard way from Moonstone to Chicago, and from there on to the musical capitals of Germany, to come back to America, successful and accomplished as one of the world famous Wagnerian sopranos of the turn of the century.^{1}

In the Preface written in 1932 for a new printing of the novel, however, Cather finds fault with its title, borrowed from a painting by Jules Breton in the Chicago Art Institute that had much impressed her at the time of her first visit to Chicago during her student years at the University of Nebraska in Lincoln. The peaceful landscapes painted by Millet, Breton, Théodore Rousseau, and the other members of the Barbizon School reminded her of her own Nebraska prairie environment; their sturdy peasant girls and women were not unlike the hard-working pioneer women she had seen toiling about the farms around the town of Red Cloud where she had arrived at nine from her native Virginia.^{2}

Jules Breton’s painting shows a young peasant girl on her way to the fields at early morning, looking up to listen to a lark. The title, Cather writes, “was meant to suggest a young girl’s awakening to something beautiful.” She fears, however that “readers take it for granted that the ‘lark song’ refers to the vocal accomplishments of the heroine,” which she deems “altogether a mistake,” for “her song,” she contends, is not “of the skylark order.”

1 The novel is divided into Parts and chapters. These will be indicated in the text of this article with the page number of the Virago Press edition (London, 1982) which follows the revised 1937 text and contains Cather’s 1932 Preface. General information about Cather is supplied by James Woodress’s biography of the novelist.

2 Cather had made the pilgrimage to Barbizon during her first journey to France, in 1902. She sent descriptions of it to the Nebraska State Journal.

ANDROGYNY IN WILLA CATHER

Was it that the lark eventually came to look of too plain a feather for the character of Thea and its song too much of the rustic kind? Readers of the novel may not think so. The diva they encounter in the last pages in the role of Siegelinde in Richard Wagner’s The Valkyrie in the full bloom of her beauty and success may have shed the fierce awkwardness of the young Moonstone girl, but some of the girl’s healthy sturdiness and earthy composure still clings to both the diva she has become and to her various operatic interpretations; and besides, what other bird could Cather have chosen to represent her? The eagle does not sing, nor does the swan (unless approaching death), and certainly Thea is no dove, that traditional ultrafeminine cooer, the bird of sacred Aphrodite/Venus whose part we read she occasionally sings alternately with the more congenial Elizabeth in Wagner’s Tannhäuser. The nightingale of course may come to mind, but apart from her connoting a lighter soprano than Thea’s rich mezzo and the coloratura she does not possess, the nightingale metaphor supposedly came too close the real Jenny Lind (“the Swedish Nightingale”) who, we know, was not Cather’s model for Thea.^{3}

Yet in view of Thea’s consistent metaphorization in the novel, the lark may serve as an appropriate substitute for the non-singing eagle. Both sun birds, soaring high into the sky, they are also male images in an homogeneous symbolical network which makes Thea appear as the total being, both male and female in its essence. The present essay’s purpose is to draw out the various representations of androgyny, and in doing so, to complete and perhaps invert/subvert the more maternal images of birth and matrilinearity set down by critics such as Ellen Moers, Judith Fryer, Josephine Donovan, or Sharon O’Brien, to name but a few.

ÉTUDES ANGLAISES

“Every artist makes himself born,” Harsanyi, the piano teacher in Chicago, tells the seventeen-year-old Thea, but he also warns her that “it is very much harder than the other time, and longer” (Part II, chap. 3, p. 221). In The Song of the Lark, the process of Thea’s (re)birth as an artist is being set off when, as a twelve-year-old adolescent in a crowded family, she is given a room of her own in the attic of the house. This little “attic room” enables her to listen to the “voice within herself” and to concentrate on her “inner self.” Very early in life, she is thus made aware that she is two persons leading a “double life” and that only her “inner voice” can restore to her a “sense of wholenesss.” The room is her “shell”; it is described as a “sunny cave,” “a little morning cave facing the sun” (II, 11, 300), but also as a snug cubbyhole drenched by moonlight during the nights. Thea’s attic room is thus the first womb image in the novel, the nest out of which the lark will be hatched.

3 Jenny Lind, however, is mentioned three times in the novel, twice together with Henrietta Sontag as the two artists who “opened a new world” to Theodore Thomas, the real-life conductor of the Chicago Symphony Orchestra, in a conversation with the fictive Harsanyi (I, 6, 260-61). The clichéed nightingale metaphor, however, is to be found in a review Cather wrote for the Nebraska State Journal of 25 January 1895 in which she stated, at age 21, that ”married nightingales seldom sing,” a bold pronouncement quite in keeping with the argument of the present article (in The Kingdom of Art). On Olive Fremstad, the Wagnerian soprano who is the model for Thea, see a recent article in the October 1997 issue of the Yale Review by Peter G.

Davis.

The voice once openly recognized during her apprenticeship in Chicago — and there are distinct stages in this recognition which are as many “revelations” on Thea’s epiphanic road to artistic achievement — puts an end to the double life Thea had been living so far and by making her “whole” also determines her rebirth to the artistic life of an opera singer.

The scene of Thea’s (re)birth is not Chicago, however, the hustling, brawling, male city of skyscrapers and meat packers, the “hog butcher” of Carl Sandburg’s poem, but the desert of Arizona and the ruins of the ancient cliff-dwelling Indians in a place Cather calls “Panther Canyon.” Ellen Moers has dubbed this landscape of tablelands and gorges “the most thoroughly elaborated female landscape in literature” (Moers 258), a statement developed to quasi-formulaic proportion by later female/feminist critics.

In light of the rebirth enacted in the “folds” of Panther Canyon, Thea’s two years in Chicago are a “wandering in the desert,” a time also compared to the period of Chaos before the moment of Creation. It is a time of doubts and frustrations, when her goal seems as far away as God to the mystic lost in acedia, when she feels submerged under what Henry James in The Portrait of the Lady has called “the bottomless idiocy of the world” (which Cather perhaps remembered when she entitled section III of the novel “Stupid Faces”).

Thea is snapped out of it, however, by the “revelation” that is visited upon her in Panther Canyon, when in another “sunny cave” high up in the San Francisco mountains, the artist is finally being born. The nineteen-year-old Thea who emerges from it has a clear image of her future priorities; she has integrated and transcended her past and glimpsed the essential one-ness and continuity of Life, Nature and Art. She no longer feels “divided”, her “two selves” have become one: Kronborg has been born.

The whole section IV of the novel, “The Ancient People,” is devoted to those momentous two months of Thea’s life. Here is a description of Panther Canyon for which Moers, taking her cue from Freud’s cautious representation of the female reproductive organs as a most “complicated topography,”^{4} would willingly have coined the term “gynetopography,” were it not prone, as she claims, to litter the “air of literary debate” with superfluous Greek words (Moers 257).”

Panther Canyon was like a thousand others — one of those abrupt fissures with which the earth of the Southwest is riddled […] it was accessible only at its head. The canyon walls, for the first two hundred feet below the surface, were perpendicular cliffs, striped with even-running strata of rock. From there on to the bottom the sides were less abrupt, were shelving, and lightly fringed with pinions and dwarf cedars. The effect was that of a gentler canyon within a wilder one. The dead city lay at the point where the perpendicular outer wall ceased and the V-shaped inner gorge began. There a stratum of rock, softer than those above, had been hollowed out by the action of time until it was like a deep groove running along the sides of the canyon. In this hollow (like a great fold in the rock) the Ancient People had built their houses of yellowish stone and mortar. The overhanging cliff above made a roof two hundred feet thick. The hard stratum below was an everlasting floor…

In both walls of the canyon the same streak of rock had been washed out, and the long horizontal groove had been built up with houses… The canyon twisted and wound like a snake, and these two streets went on for four miles or more, interrupted by the abrupt turnings of the gorge, but beginning again within each turn… (IV, 2, 369-70; the various womb images have been italicized).

Further still, the canyon is “this crack in the world,” “an open gulf,” a “chasm,” a “cleft in the heart of the world”.

In the middle of this “female” landscape, in the snake-like groove carved out of the rock, are the dwellings of the Ancient People, the ancestors of the present Pueblo Indians who at the time of the writing of The Song of the Lark were not yet known by their present name of Anasazi (a Navajo term meaning “ancient enemies”). Established from prehistoric time in the area contiguous to the “four corners states” of Colorado, Utah, Arizona, and New Mexico, the Anasazi by successive waves along the centuries had been driven out of their canyons, the last ones leaving well before the “discovery” of America by Columbus. It was thought for a time that they were killed off in their incessant warfare with the Plain Indians, but it is now established that they left after protracted periods of severe drought to settle along the Rio Grande where they can still be found.

There Thea appropriates one of the Indians’ “rock-rooms.” It is another “sunny cave,” another “nest in a high cliff, full of sun” (IV, 2, 371). In it she feels secure and in perfect symbiosis with her environment, like an unborn child “swimming in the liquid gold of the sun.” She does not think or worry, she becomes sheer sensation, her body converted into “a mere receptacle for heat,” becoming … a colour… a continuous repetition of sound” (IV, 2, 373).

4 In his “Tenth Introductory Lecture to Psycho-analysis,” cited by Moers (254).

There also she experiences the temptation that was visited upon many a mystic or a mythological hero before her — and who knows? the unborn child — the temptation to dream her life out in the blissful euphoria of some womb-like spot, away from life and its entanglements. Erich Neumann has called this a state of “uroboric incest,” a (self-)gratifying lethargy akin to death, and full of perils for the weak-hearted individual.

The “revelation” granted to Thea in her little rock-room is the major epiphany of the novel; it is the sudden intuition that grasps her of the true meaning of the Ancient Indian women’s art of pottery whose sherds still litter the floors, and of its relation to her own. The Indian women’s precious decorated bowls and jars, she understands in a flash of illumination, are but “the envelope and sheath” of water without which there is no life, just as her throat is the vessel of her voice which, being breath, is that other precious element without which life could not be, air.

The stream and the broken pottery: what was any art but an effort to make a sheath, a mould in which to emprison for a moment the shining, elusive element which is life itself -life hurrying past us and running away, too strong to stop, to sweet to lose? The Indian women had held it in their jars. In the sculptures she had seen in the Art Institute, it had been caught in a flash of arrested motion. In singing, one made a vessel of one’s throat and nostrils and held it on one’s breath, caught the stream in a scale of natural intervals” (IV, 3, 378; emphasis added).

Decorated pots and bowls to contain water, the singer’s throat to sheathe her voice which is air, the fire to harden the clay and steel an artist’s temperament, produce Art which is Life itself, a moment snatched off our common mortality. Art is “the seed” created by the conjunction of female (the sheath, vessel, receptacle) and male (air, fire, and for Cather, water),^{5} the necessary combination of womb and semen for a life to be born.

Thea contains both male and female elements in herself, the sheath^{6} of her throat, larynx… the physical apparatus of her singing, and her beautiful, natural voice, a voice she had always dimly known she possessed. Her Moonstone teacher had discerned it (“a nature voice… breathed from the creature and apart from language, like the sound of the wind in the trees, or the murmur of water,” I, 11, 97)), Harsanyi stood in awe of it when he discovered it in Chicago (see the end of chap. 3, III, 237-38). With her final revelation in Panther Canyon Thea is indeed made “whole,” the entire creature contained within its androgynous self.

5 I am perfectly aware that water is more universally an element related to the moon, and thus to femaleness, but in its constant comparison to air, it seems Cather makes it a male element. Water, the undifferentiated matter of the Origins in many cosmologies, may also be seen in its male avatar of the Heavenly rain fecundating the Earth in the first act of Creation, similar to the Breath or Spirit of God in the Book of Genesis in the Bible. Yahvé, for instance, is compared to rain (Hosea: 6, 3), to dew (idem: 14, 6); in Greek mythology, to give but one example, Zeus impregnates Danae through a golden shower. It is this male aspect of fecundating semen bursting forth and creating life in the womb that Cather seems to have privileged .

6 Ellen Moers very appropriately reminds us of the Latin meaning of vagina, which is but the sheath or scabbard of the male organ (256).

For the womb-like landscape of Panther Canyon to produce life, to deliver a new Thea, as, “far back in the night of the past,” it had produced the art of the Ancient Indians, it must needs contain the many fecundating male elements that Moers and the feminist critics in her wake have failed to notice. These stand out in the blaring sun- and moonlight, inescapable for who reads Cather’s description of the canyon landscape with an alerted eye. Some of these elements in turn symbolize the intruding maleness of Fred Ottenburg, Thea’s patron, short-time lover and later-life husband whose ever-faithful presence at the side of the self-sufficient androgynous Thea will make of his life a “heroic discipline” (as he tells her himself when the reader meets him again some ten years later. “But loving you is a heroic discipline. It wears a man out” (VI, 10, 559)).

In Panther Canyon, however, Fred undoubtedly is Thea’s sexual awakener, his masculinity represented by all the phallic elements that stand up in an aggressive way in the bowl-like landscape of the canyon:

the “clump of sharp bayonet leaves” of the yuccas, out of which rises “a tall stalk hung with greenish white bells with thick, fleshy petals” (IV, 5, 383), an adequate representation, if any, of male genitalia;

the “niggerhead cactus… thrusting its crimson blooms out of every crevice in the rock” (idem);

the jutting rocks, the tall pine trees on top of the mesa, the cedars and pinions that stand “dark and rigid, like bronze” (7-401) on the slopes and assert their presence in the “coppery fire” of the sun;

and, of course, the hard blazing sun itself, the pervasive impregnator, and the spermatic fluid of the stream winding its way on the floor of the canyon, compared all through Section IV to the air between the canyon walls, the “blue-air river,” the “open gulf” in which “arrow-shaped” swallows “swim all day long.”

Thea’s fierce awkward womanhood is aroused by Fred, the “savage blonde” (as one of her fellow students called her in Chicago) is awakened to the stirrings of sexual desire. She “delights” in “the pleasant vigour of his person,” she feels an invigorating “life in the air… something coming and going, a rhythm of feeling and action” (6, 392). “Never,” the reader is told, had she “felt this pleasant excitement about any man before” (7, 399). And the perceptive Fred pays homage to her new loveliness and feminity.

There is no lovemaking, however, in the expected sense. Thea and Fred spend their days horse-riding on top of the mesa, climbing up and down the canyon slopes, challenging each other at throwing rocks into the open gulf of the canyon, fencing with single-sticks in lieu of foils, untiringly exploring the rock houses in their winding groove on both sides of the canyon. From afar they “look[ ] like two boys” (5, 384). When they kiss, since kiss they eventually must, Thea does not hide her face on Fred’s shoulder, as any maiden was expected to do, “she [rises] a little on her toes, and [stands] straight and free”; she “[rises] to meet him” and “[kisses] him without constraint and embarrassement” (7, 404-405). With the same rising movement, she “[springs]” to her feet and “[stands] rigid” on the edge of the rock to salute a golden eagle that comes “sailing over the cleft in which she lay”, “full of warmth, lassitude and physical content” as after an embrace (6, 398).^{7}

The reborn Thea is no maiden awaiting a man to be furnished with a destiny. As her father already knew in Moonstone, she is not “the marrying kind” and for Fred certainly she is no “nest-building bird” either. He won’t propose to her in the traditional way, offer her “a comfortable flat in Chicago, a summer camp up in the woods, musical evenings, and a family to bring up” (6, 393-94). He merely promises to himself to “keep an eye on her” (idem), “make it his business in life to get her on” (7, 413),^{8} well aware of, and acquiescent to, being “an instrument” and not the worshipped “Apollo of a homey flat” (6, 391). And in still another reversal of the traditional betrothal scene, it is Fred and not the tremulous future bride who asks the ritual question: “Are you going to play fair?” hoping she “will want him around” for better reasons than “to play with” (7, 403).

That marriage, as Fred perceives, would never be an “end” for Thea but merely an “incident” is predicated by the various male/female images found in the novel which, singly or compounded, underpin Thea’s essential androgyny.

Thea’s voice (a male element similar to breath and air as we have seen above) is like “a wild bird that had flown into his studio,” the first time Harsanyi hears her sing (II, 3, 232). It is the great tawny eagle that comes sailing over the rim of Panther Canyon, which is both her own striving self and the proposing Fred and which Thea salutes as a symbol of “endeavour, achievement, desire, glorious striving of human art” (IV, 6, 399). The lark is a more modest achiever, but soaring high into the sky from the flat lands Thea is crossing on her way from Chicago to Moonstone, it “sings a new song” on a “young and fresh and kindly” land where immigrants “from old, sad countries [are] given another chance” (II, 8, 276-277). An appropriate symbol for the daughter of Norwegian-Swedish immigrants who “makes good” on American soil, the lark, like the eagle, is a solar image, but nesting as it does in the rich furrows of the bountiful earth, it associates sky, sun and earth in yet another image of androgyny. Such are also the “arrow-shaped” swallows that cruise the blue air gulf between the Canyon walls, which combine the male graphic of their shape with the maternal image of their “nest-building” capacity (explicitly compared in the text with the ancient cliff-dwellers’ (IV, 3, 375). The arrow, besides, connotes the undeviating determination of Thea’s dedication to a career out of the ordinary, felt by all ever since she was “two feet high,” and similarly metaphorized as the trajectory of a “rifle ball” by Fred. Eagle, lark, arrow-shaped swallow, anything but the dove cooing her sentimental stuff in the traditional dovecotes of femaleness.

7 The very first kissing episode is rather a fight, a taming of wild horses, reminding one of one of Thea’s mythological etymologies: the daughter of the centaur Chiron, transformed into a mare by Poseidon and later into the Horse constellation (J. E.

Zimmermann. Dictionary of Classical Mythology. New York: Bantam, 1964; 262).

8 Fred Ottenburg is a rich brewer’s son, well introduced in the world of music and culture in both Europe and America.

At one of their meetings during the time Cather was composing her profile of “Three American Singers” for the December 1913 issue of McClure’s, Olive Fremstad had told Cather a piece of sober truth that she would remember when she wrote The Song of the Lark. Fremstad had said: “We are born alone, we make our way alone, we die alone.” This of course may be true for anybody, and is not entirely so for Thea who has never lacked dedicated friends and supporters all through her years of apprenticeship and after. But it characterizes the fundamental loneliness of the artist vis-à-vis her art and its exacting demands. The diva Thea turns into has not much of a personal life indeed, not to mention a normal married life and children. The images for this state of proud loneliness in the novel are the great pine trees that dot the Navajo tableland of the Panther Canyon area at the beginning of “The Ancient People” section.

The great pines stand at a considerable distance from each other. Each tree grows alone, murmurs alone, thinks alone. They do not intrude upon each other. […] Each tree has its exalted power to bear.” (IV, 1, 367)

The Ur-androgyny at the outset of many mythical cosmogonies where it joins Heaven and Earth together to initiate the very first act of Creation, the tree combines the male erectness of its shape and the female abundance of its fruit. The pine, moreover, was the tree dedicated to Cybele, the Great Phrygian Mother, and it was Attis himself, her slain youthful lover. And male/female are likewise the lesser trees that cling to the slopes of the canyon, the wild pinions, cedars and chokecherries that embody Thea’s obstinate will and endurance, the cottonwood seedlings that shade her ritualistic bathing in the canyon stream.

At the close of the novel, when the readers leave Thea during a performance of The Valkyrie at the Met, they behold a beautiful soprano who has come into “full possession of things she had been refining and perfecting for so long” (VI, 11, 571). Usurping her character’s consciousness, Cather leaves her “feel[ing] like a tree bursting into bloom” (idem), in a powerful image joining both male and female in the orgasmic celebration of life (and art).

It may not be fortuitious either if the diegesis proper of the novel (barring the short proleptic Epilogue that sums up Thea’s career for another eight years from the vantage point of Moonstone and her aunt Tilly) closes on Thea singing Siegelinde in that very opera. Siegmund and Siegelinde, the Volsung pair, the twin brother and sister of Germanic mythology, are, like so many twins in mythology, the two faces of the same coin of androgyny, “the identity-in-essence of the separate-in-form,” as Southern writer Andrew Lytle has called the brother/sister pair in a novel much concerned with the striving towards wholeness.

Thea certainly is no “maiden flower” as her Moonstone teacher perceives it so early in her life, no lissom seductress of men like the Kundry in Wagner’s Parsifal which will undoubtedly be part of her operatic repertoire. To Professor Wunsch she appears like “a flower full of sun… like the prickly-pear blossoms that open there in the desert; thorn[y] and sturd[y], but wonderful” (I, 13, 122).

The sunflower image by its constant association in the novel with the moonflowers of Thea’s Mexican neighbours’ garden echoes the pervasive sun/moon symbolism of The Song. It overlays all the individual male/female images that shape Thea’s identity.

“Moonstone,” her native Colorado town, is a “desert town,” scorched by the sun and surrounded by stretches of hard glittering sand dunes that turn white or lavender in the summer nights of full moon. These will for ever make up the inner landscape of her soul, as does Panther Canyon “swimming in the liquid gold” of the blazing sun.

As much, however, as Thea appears a child of the sun, throbbing with solar life and energy, as much do we see her “pulsing with ardour and anticipation” in her attic room at night, drenched by moonlight. The restless adolescent roams the moonlit streets of her town, watches wild rabbits (an archetypal model of fertility for which she can muster but tepid enthusiasm) cavorting under the full moon and decides to live only for “impossible” things. Braving small town propriety, she spends nights singing and dancing in nearby Mexican Town, when “the moon itself looked like a great pale flower in the sky” (II, 10, 291). Later, she will remember the moonflowers of Mrs Tellamantez’s garden, the sand ridges turning lavender in the light of the moon and these too will overflow into her music and become part of her artistic self.

We may conclude this tentative essay on the various representations of androgyny in The Song by going back to the figure of Siegelinde in the last chapter of the novel, Siegelinde at last reunited with Siegmund and lost in a swooning ecstasy of love during a night so soaked in moonlight that they think a sunny spring day has dawned upon them. Or rather, we might conclude with the “seed” of that precious night and with Siegfried evoke that other “savage blonde” of Germanic mythology, both heroic male and submissive female, the Valkyrie Brunhild herself. We do not actually see Thea sing the role, but all of Thea’s career leads to it, the crowning role of all Wagnerian sopranos, and so we close the novel not so much in suspense as in expectation of it.

With the figure of the Valkyrie, we leave Thea as she has appeared all through her story of apprenticeship and success, “a personality… larger than we are accustomed to see them… that carrie[s] across big spaces and expound[s] among big things,” as she is seen by Fred in Panther Canyon (IV, 6, 397). And, in the full light of her Greek name, remember, as Cather well knew,^{9} that Hesiod in his Theogony made her a Titan who, forced by her brother Hyperion, became the mother of golden Helios and fair Selene, and also of rosy Dawn where Sun and Moon ever so briefly come together.

If a voice “simply is the mind and is the heart,” as Cather has Fred Ottenburg declare near the end of the novel (VI, 5, 509), then Thea, both male intellect and female sensibility, truly appears as that self-contained mythic creature, the “imperishable daughter of music” of Cather 1932 Preface.

Anne FOATA Université Marc Bloch, Strasbourg

Note added in proof. On the subject of androgyny in general cf. the book by Marie-Claude Perrin-Chenour in the collection “Voix Américaines” published by Belin in 1998 (especially chap. 3) which came to my notice after submitting the present paper.

9 Cather had studied Latin and Greek with a Red Cloud neighbour but had to take a whole year of cramming in those two languages in order to be allowed to enter the University of Nebraska where they were part of her four-year cursus (Woodress, 43, 53, 71-72).

BIBLIOGRAPHY

Anasazi (The). Books and brochures collected at the various cliff-dwelling sites of the Southwest; e.g. Richard J. Ambler. The Anasazi. Flagstaff, Arizona: Museum on North Arizona Press (no date given).

Cather, Willa. The Song of the Lark. Boston: Houghton Mifflin, 1915. Revised edition, 1932. Reprint by the University of Nebraska Press (Lincoln) in 1978.

The Kingdom of Art. Willa Cather’s First Principles and

Critical Statements 1893-96. Ed. Bernice Slote. Lincoln: The University of Nebraska Press, 1966; p. 175-76.

Donovan, Josephine. After the Fall. The Demeter-Persephone Myth in Cather, Wharton, and Glasgow. University Park: Pennsylvania State University Press, 1989; p. 85-127.

Eliade, Mircea. Traité d’Histoire des Religions. Paris: Payot, 1964 (revised

edition, 1970). Fryer, Judith. Felicitous Space. The Imaginative Structures of Edith Whar-

ton and Willa Cather. Chapel Hill: the University of North Carolina Press, 1986; p. 232-40 & 293-300.

Graves, Robert. The Greek Myths, rev. ed. Penguin, 1960. Hesiod. Theogony. Translated into English by Dorothea Wender. Har-

mondsworth, Penguin, 1976; v. 134-38 & 373-76. Jung, Carl G. Psychologie et Alchimie. French ed.; Paris: Buchet/Chastel,

1970.

Mysterium Conjunctionis (2 vol.). French ed.; Paris: Albin

Michel, 1980. Lytle, Andrew. The Velvet Horn. New York: McDowell, Obolensky, 1957;

p. 111. Moers, Ellen. Literary Women. New York: Doubleday, 1976; p. 257-63 in

the 1985 Oxford University Press (New York) reprint. Neumann, Erich. The Origins and History of Consciousness. Bollingen

Series XLII. Princeton University Press, 1970; p. 260-312.

The Great Mother. Bollingen Series XLVII. Princeton Uni-

versity Press, 1972. O’Brien, Sharon. Willa Cather. The Emerging Voice. New York: Oxford

University Press, 1987; p. 170-73 & 403-20 of the 1988 Random House (Ballantine) reprint.

Woodress. James. Willa Cather. A Literary Life. Lincoln: University of Nebraska Press, 1987.